OpenStudy (anonymous):

GIVING MEDALS in right triangle ABC, A=76 degrees, a=13 an

OpenStudy (anonymous):

for the angles, you know A= 76, C = 90, that would leave B= 14 i found this with the angle sum, all angles of a triangle add to be 180. so 76 + 90 + B = 180 166+ B = 180 B = 14 now for the angles, use the law of sines. which is (sin A)/(a) = (sin B)/(b) = (sin C)/(c) since you know angle A and side a and Angle B, find side b (sin A)/(a) = (sin B)/(b) (sin 76)/(13) = (sin 14)/(b) b*sin 76 = 13*sin 14 b = (13*sin 14)/(sin 76) b = 3.24 use Pythagorean theorem to solve for side c a^2 + b^2 = c^2 13^2 + 3.24^2 = c^2 179.50 = c^2 √179.50 = c 13.40 = c Follow me on www.twitter.com/devanshgaur